OpenStudy (anonymous):

write an equation in slope intercept form of the line whose y-intercept is -3 and the line is parallel to the line whose equation is 3x-2y=9

OpenStudy (anonymous):

-2y=-3x+9 y=3/2x=9 so y=3/2x-3

OpenStudy (anonymous):

@mit........ pls elaborate

OpenStudy (marco26):

y=mx+b Since parallel lines have the same slope, the required equation of the line has the same slope as that of 3x-2y=9 Now we know all the values, plug them in and that's the answer

OpenStudy (anonymous):

okay tahnks

OpenStudy (anonymous):

thanks